OpenStudy (anonymous):

the area of a square is numerically equal to five times its perimeter find the length of a side of the square

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Can you explain to me how you got that answer?

OpenStudy (anonymous):

sheg is wrong.

OpenStudy (anonymous):

I'm really horrible at math

OpenStudy (anonymous):

Never mind he is not lol, but there is other possibilities.

OpenStudy (anonymous):

sheg is right a^2=5p 4a=p a^2=5*4a a^2=20a a=20

OpenStudy (anonymous):

genius is it wrong !!!!!!

OpenStudy (anonymous):

genius i m asking to u ??????????

OpenStudy (anonymous):

No you are not wrong, I already said that anyway.

OpenStudy (anonymous):

If P or perimieter = 4s, s is side. And area=5p, then, we can put this in to an equation like this: 20 * 4s = p. p = 80 a = 400